A wire of length πŸ‘πŸ“πŸŽ π’„π’Ž and radius 𝟏. πŸ“ π’Žπ’Ž is stretched by load of 𝟏𝟎 π’Œπ’ˆ and increased by 𝟐. πŸ“% of the original length. Find the stress, strain and Young’s modulus of the material of the wire


1
Expert's answer
2021-11-29T11:43:51-0500

We can find the stress as follows:


"\sigma=\dfrac{F}{A}=\dfrac{mg}{\pi r^2},""\sigma=\dfrac{10\ kg\times9.8\ \dfrac{m}{s^2}}{\pi\times(1.5\times10^{-3}\ m)^2}=1.38\times10^7\ \dfrac{N}{m^2}."

We can find the strain as follows:


"\epsilon=\dfrac{\Delta l}{l}=\dfrac{3.5\ m\times0.025}{3.5\ m}=0.025."

We can find the Young’s modulus of the material of the wire as follows:


"E=\dfrac{\sigma}{\epsilon}=\dfrac{1.38\times10^7\ \dfrac{N}{m^2}}{0.025}=5.52\times10^8\ \dfrac{N}{m^2}."